OpenStudy (anonymous):

is this the right answer this (7 x 12) + 4(-8 x 4) = -44

OpenStudy (anonymous):

(7 x 2) + 4(-8 x 4) = -44 14 + 4 (-32) = -44 14 - 128 != - 44 not right

OpenStudy (goformit100):

its correct

OpenStudy (kropot72):

@nanananabatman Yes. -44 is the correct answer.

OpenStudy (anonymous):

OHHhh.. my bad. sorry :((